# A Note on Two Useful Lemmas in Complex Algebraic Geometry

### 16 Apr 2026 v2

@@ Abstract
We establish two useful lemmas in complex geometry. First, we investigate the homological properties of differential operators intertwined by the Dolbeault Laplacian on compact Hermitian manifolds. By utilizing the mapping cone of Koszul complexes, we prove that the analytical obstructions to solving specific overdetermined PDE systems reduce entirely to the finite-dimensional Dolbeault cohomology. Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
Keywords: Dolbeault cohomology, Koszul complex, Complex manifold, Hyperbolicity
@@ 1 Introduction and Background
In accordance with the spirit of rigidity exemplified by Hartogs’ theorem, this paper investigates how rigid global structures in complex geometry can be enforced by partial algebraic or homological constraints. We divide our exploration into two distinct but philosophically related results.
In Section 2, we turn our attention to differential operators on compact Hermitian manifolds. In the study of integrable systems, operators intertwined by a Laplacian exhibit highly constrained solution spaces. We abstract this phenomenon using homological algebra. By establishing an acyclicity lemma for intertwining chain complexes, we demonstrate how the Dolbeault Laplacian acts as an invertible intertwining operator on the space of exact forms. This allows us to prove that the infinite-dimensional analytical obstructions to solving certain coupled PDE systems collapse entirely into the finite-dimensional Dolbeault cohomology.
In Section 3, we transition from the rigidity of differential equations to the rigidity of mappings. While Hartogs’ theorem relies on boundary integrals within bounded domains, transitioning to compact projective varieties requires global algebraic geometry. We present a rigidity lemma demonstrating how continuous, fiber-wise analytic mappings can be automatically elevated to global bi-holomorphic isomorphisms when anchored by a transverse algebraic intersection, specifically a very ample hypersurface, on a Kobayashi hyperbolic manifold.
2 A Lemma on Holomorphic Differential Equations
We begin our investigation by formulating a purely algebraic statement. By localizing the homological complexity to a specific submodule where an intertwining operator acts as an isomorphism, we equate the cohomology of the total space with that of the quotient.
@@ 2 A Lemma on Holomorphic Differential Equations
Let $0\rightarrow A\rightarrow B\rightarrow C\rightarrow 0$ be a short exact sequence of $R$-modules, and let $f,g,h\in\operatorname{End}_{R}(B)$ satisfy $fg=gh$, so that the sequence $K(B)$ of $R$-modules given by
[…]
is a chain complex. Suppose that $h(A)\subseteq g(A)=A$ and $g|_{A}$ is a monomorphism. Then $0\rightarrow A\rightarrow A^{\oplus 2}\rightarrow A\rightarrow 0$ forms an acyclic subcomplex $K(A)$ of $K(B)$, such that the cohomology groups of $K(C):=K(B)/K(A)$ are isomorphic to that of $K(B)$.
@@
Since $fg=gh$, we deduce that $f(A)\subseteq A$. Thus, the endomorphisms $f,g,h$ restrict to the submodule $A$ and descend to the quotient $C=B/A$. We therefore obtain a short exact sequence
$0\longrightarrow K(A)\longrightarrow K(B)\longrightarrow K(C)\longrightarrow 0$
of complexes, which induces a long exact sequence
$\dots\longrightarrow H^{i}(K(A))\longrightarrow H^{i}(K(B))\longrightarrow H^{i}(K(C))\longrightarrow H^{i+1}(K(A))\longrightarrow\cdots$ (1)
of cohomology groups.
We claim that the subcomplex $K(A)$ is acyclic. To verify exactness at the first term, we take $a\in A$ satisfying g(a)=0-g(a)=0. Since $g|_{A}$ is a monomorphism, we obtain $a=0$. Next, to establish exactness at the middle term, we take $(a,b)\in A\oplus A$ such that $f(a)+g(b)=0$. Because $g(A)=A$ and $g|_{A}$ is injective, in fact $g|_{A}$ is an automorphism. Consequently, there exists a unique $c\in A$ such that g(c)=a-g(c)=a. Applying $g$ to $h(c)$ yields $g(h(c))=(gh)(c)$, which becomes f(g(c))=f(a)=f(a)f(g(c))=f(-a)=-f(a) by virtue of the intertwining relation $fg=gh$. Since $f(a)+g(b)=0$, we have g(b)=f(a)=g(h(c))g(b)=-f(a)=g(h(c)). Because $g|_{A}$ is injective, we conclude $b=h(c)$. Thus, it holds that (a,b)=(g(c),h(c))(a,b)=(-g(c),h(c)), proving exactness at $A\oplus A$. Finally, to prove exactness at the terminal term, we choose an arbitrary $c\in A$. Since $g|_{A}$ is surjective, there exists $b\in A$ such that $g(b)=c$, demonstrating surjectivity. The acyclicity of $K(A)$ is therefore established.
The exactness of sequence (1) then forces $H^{i}(K(B))=H^{i}(K(C))$ for all $i\in\mathbb{Z}$. ∎
We now translate this abstract homological machinery into the geometric language of differential equations on compact Hermitian manifolds. By choosing our short exact sequence to reflect the Hodge decomposition of differential forms, we let the Dolbeault Laplacian assume the role of the intertwining operator. Because the Laplacian acts invertibly on the orthogonal complement of the harmonic forms, Lemma 2.1 establishes a dimension reduction: the defects of operators are completely captured by the finite-dimensional Dolbeault cohomology.
@@
Let $X$ be a compact Hermitian manifold, and $L,A$ linear differential operators on $E:=\Omega^{p,q}(X)\cap\operatorname{Ker}(\bar{\partial})$. Suppose that $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$. Then, the following properties hold:
  1. 1.
    a solution $\omega\in E$ to $L\omega=\Delta_{\bar{\partial}}\omega=0$ is precisely a harmonic form of which Dolbeault cohomology class lies in the kernel of the induced operator $L^{*}\colon H^{p,q}(X)\rightarrow H^{p,q}(X)$;
  2. 2.
    for any $\alpha,\beta\in E$ satisfying $A\alpha=\Delta_{\bar{\partial}}\beta$, there exists $\omega\in E$ simultaneously solving $\Delta_{\bar{\partial}}\omega=\alpha$ and $L\omega=\beta$ if and only if the cohomology class $\alpha]=0$ and $\beta]\in\operatorname{Im}(L^{*})$.
@@
Consider the short exact sequence of vector spaces
$0\longrightarrow\operatorname{Im}(\bar{\partial})\cap E\longrightarrow E\longrightarrow H^{p,q}(X)\longrightarrow 0.$
Since the Dolbeault Laplacian $\Delta_{\bar{\partial}}$ commutes with $\bar{\partial}$, it restricts to an endomorphism on $E$. By Hodge theory, we have $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$. Furthermore, as Green’s operator commutes with $\bar{\partial}$, the restriction of $\Delta_{\bar{\partial}}$ to the subspace $\operatorname{Im}(\bar{\partial})\cap E$ is an automorphism.
Moreover, for any $\alpha\in\operatorname{Im}(\bar{\partial})\cap E$, there exists $\omega\in\operatorname{Im}(\bar{\partial})\cap E$ such that $\alpha=\Delta_{\bar{\partial}}\omega$. Therefore, we obtain that
$A\alpha=A\Delta_{\bar{\partial}}\omega=\Delta_{\bar{\partial}}(L\omega).$
Since $L\omega\in E$, we have $\Delta_{\bar{\partial}}(L\omega)\in\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, which guarantees that the operator $A$ also preserves the subspace $\operatorname{Im}(\bar{\partial})\cap E$.
The hypotheses $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$ thus ensure that the operators under consideration satisfy all premises of Lemma 2.1. Using the notations in the proof of Lemma 2.1, we have the isomorphism $H^{i}(K(E))=H^{i}(K(H^{p,q}(X)))$ for all $i\in\mathbb{Z}$. Because $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, the induced operator of $\Delta_{\bar{\partial}}$ on the quotient $H^{p,q}(X)$ is identically zero. The complex $K(H^{p,q}(X))$ thus simplifies to
[…]
Now, straightforward computation yields $H^{0}(K(H^{p,q}(X)))=\operatorname{Ker}(L^{*})$ and
$H^{1}(K(H^{p,q}(X)))=\operatorname{Ker}(A^{*})\oplus\operatorname{Coker}(L^{*}),$
which immediately translate to the two desired properties. ∎
3 A Lemma on Kobayashi Hyperbolic Manifolds
@@
Step 1: Since $H\in\operatorname{div}(X)$ is ample, the complex manifold $X$ is projective by Kodaira’s embedding theorem. Let
@@
be the punctured Riemann sphere consisting of the regular values of $f$. For any $t\in\Sigma$, the fiber $X^{t}:=f^{-1}(t)$ is a smooth projective subvariety of $X$ of pure dimension n1n-1. Since $\phi^{t}:=\phi|_{X^{t}}$ is holomorphic and $Y$ is projective, Remmert’s proper mapping theorem and Chow’s theorem ensure that the image $Y^{t}:=\phi(X^{t})$ is a projective subvariety of $Y$. We first establish the injectivity of the restriction $\phi^{t}$ for any fixed $t\in\Sigma$. For the sake of contradiction, we now assume that $\phi^{t}$ is not injective.
@@
Suppose now that $\dim(Y^{t})=n-2$. Take a generic point $y\in Y^{t}$. Then the fiber $C$ of $\phi^{t}$ over $y$ is a projective curve. Since $H\in\operatorname{div}(X)$ is a very ample divisor, $H^{t}:=X^{t}\cap H$ is very ample in $X^{t}$, making $H^{t}\cap C$ non-empty and $\phi(H^{t})=Y^{t}$. Because $\phi|_{H}$ is injective by assumption, the mapping degree of $\phi^{t}|_{H^{t}}$ onto its image is exactly 1. By choosing $y$ to be generic, we ensure that the restriction $\phi^{t}|_{H^{t}}$ is unramified at its unique preimage $x\in H^{t}$, rendering its differential an isomorphism at $x$. As a result, the tangent space $T_{x}H^{t}$ intersects trivially with the kernel of the differential of $\phi^{t}$ at $x$, which is precisely $T_{x}C$. This proves that $C$ intersects $H^{t}$ transversally, and hence their intersection product is $H^{t}\cdot C=1$. Because $H$ is a very ample divisor, intersecting a projective curve $C$ at a single point with a multiplicity of exactly 1 forces $C$ to be rational, that is, bi-holomorphic to $\mathbb{P}^{1}$. However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem on hyperbolicity. This is a contradiction. Therefore, we have in fact $\dim(Y^{t})=n-1$, so the regular mapping $\phi^{t}\colon X^{t}\rightarrow Y$ sends its domain of definition onto a projective variety $Y^{t}$ of equal dimension n1n-1.
@@
Because $\phi^{t}_{2}$ is a surjective morphism with connected fibers between projective varieties of the same dimension, its generic fiber is a single point, implying that $\phi^{t}_{2}$ is a birational morphism. Assume to the contrary that $\phi^{t}$ admits a fiber $F$ of positive dimension. Since $\phi^{t}_{1}$ is finite, this fiber $F$ must be contracted to a point entirely by $\phi^{t}_{2}$. Therefore, the exceptional locus of the birational morphism $\phi^{t}_{2}$ is non-empty. Now, recall from birational geometry that the exceptional locus of a proper birational morphism from a smooth projective variety is uniruled, meaning it is covered by rational curves. However, as $X$ is Kobayashi hyperbolic, by Brody’s theorem, it contains no rational curves, which leads to a contradiction. This proves that $\phi^{t}$ is indeed a finite morphism.
@@
for all $t\in\mathbb{P}^{1}$. Since $Y$ is a smooth projective variety and $g\colon Y\rightarrow\mathbb{P}^{1}$ is a holomorphic fibration, Bertini’s theorem ensures that the general fiber of $g$ is smooth, and therefore normal. Thus, for a generic $t\in\Sigma$, the finite morphism $\phi^{t}\colon X^{t}\rightarrow Y^{t}$ of degree 1 onto a normal variety is an isomorphism by Zariski’s main theorem. This in particular proves that $\phi$ is injective almost everywhere.
@@
of the branched covering $\Phi|_{Z}$. Since $\Phi|_{Z}$ admits a continuous section over $\Sigma$, it is in fact not branched over $\Sigma$. Moreover, any continuous section of an unbranched analytic covering is necessarily holomorphic, so $\phi^{*}$ is holomorphic on $\Sigma$. Because $\phi$ is holomorphic along fibers and the section $\phi^{*}$ is also holomorphic, Osgood’s lemma on separate holomorphicity guarantees that $\phi$ is holomorphic on $f^{-1}(\Sigma)$. Since $\phi$ is continuous and the singular locus $X-f^{-1}(\Sigma)$ forms a proper algebraic subset, Riemann’s extension theorem guarantees that $\phi$ is holomorphic on the entirety of $X$.
Having established that $\phi\colon X\rightarrow Y$ is a proper holomorphic mapping of degree 1 between smooth projective varieties, it remains to show it is an isomorphism. But, if $\phi$ were not an isomorphism, as it is injective almost everywhere, it would be a birational map contracting an exceptional divisor to a lower-dimensional subvariety. Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$. ∎

### 22 Apr 2026 v3

# On Holomorphicity of Fiberwise Analytic Mappings

@@ Abstract
We establish two rigidity results for mappings in complex geometry. First, we investigate the elevation of continuous mappings to holomorphic ones in the setting of proper fibrations. By leveraging relative morphism spaces and the Noetherian property of their local rings, we prove that a continuous, fiber-wise holomorphic map uniquely extends to a global holomorphic map, provided it is anchored by a sequence of holomorphic sections with mutually disjoint images. This result is then applied to establish the rigidity of homomorphisms between elliptic fibrations and Abelian schemes. Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
Keywords: Complex manifold, Holomorphic mapping, Relative morphism space, Elliptic fibration, Kobayashi hyperbolicity, Birational geometry
@@ 1 Introduction and Background
In accordance with the spirit of rigidity exemplified by Hartogs’ theorem, this paper investigates how global structures in complex geometry can be enforced by partial algebraic constraints. We divide our exploration into two distinct but related results concerning the rigidity of mappings. In both settings, we demonstrate that the presence of a positive algebraic object, specifically an infinite sequence of disjoint sections or a very ample divisor, acts as a geometric anchor. When combined with appropriate local analytic conditions, this anchor elevates partial data to global holomorphic maps or global bi-holomorphic isomorphisms.
In Section 2, we turn our attention to the global extension of fiber-wise holomorphic mappings. While Hartogs’ theorem guarantees the extension of holomorphic functions across boundaries, elevating a continuous, fiber-wise holomorphic mapping to a globally holomorphic one on a proper fibration requires navigating the global geometry of the ambient space. We demonstrate that by anchoring the mapping to a sequence of holomorphic sections with mutually disjoint images, one can exploit the Noetherian nature of the local rings of the relative morphism space. This algebraic stabilization forces the mapping to be globally holomorphic, a rigidity result we apply to homomorphisms between elliptic fibrations and Abelian schemes.
In Section 3, we transition to the rigidity of proper holomorphic mappings. We present a rigidity theorem demonstrating how continuous, fiber-wise analytic mappings can be automatically elevated to global bi-holomorphic isomorphisms when anchored by a transverse algebraic intersection, specifically a very ample hypersurface, on a Kobayashi hyperbolic manifold.
2 Propagation of Holomorphicity on Proper Fibrations
We begin by investigating the conditions under which a continuous mapping, known to be holomorphic on the individual fibers of a proper fibration, can be elevated to a global holomorphic mapping. The strategy relies on translating the pointwise agreement of sections into an infinite intersection of complex analytic subsets within the relative morphism space. By invoking the Noetherian property of analytic local rings, this descending chain must stabilize, enforcing global holomorphicity.
@@ 2 Propagation of Holomorphicity on Proper Fibrations
Let $X,Y,Z$ be complex manifolds with $\dim(Y)=\dim(X)+1$ and $\phi\colon Y\rightarrow X$ a surjective proper holomorphic map with connected fibers. Let $\{s_{i}\}_{i\geq 1}$ be a sequence of holomorphic sections of $\phi$ with mutually disjoint images, and let $f\colon Y\rightarrow Z$ be a continuous mapping. Assume that $f|_{\phi^{-1}(x)}$ and $f\circ s_{i}$ are holomorphic for every $x\in X$ and all $i\geq 1$. Then $f$ is holomorphic.
@@
Let $U$ be a Zariski dense open subset of $X$ consisting of regular values of $\phi$. For any $x\in U$, the fiber $Y_{x}:=\phi^{-1}(x)$ is a compact Riemann surface. Since $Y_{x}$ is compact, the infinite set of points $\{s_{i}(x)\}_{i\geq 1}$ possesses an accumulation point in $Y_{x}$. By the analytic continuation principle for complex manifolds, any two holomorphic mappings from $Y_{x}$ to $Z$ that coincide on $\{s_{i}(x)\}_{i\geq 1}$ are identical.
Consider the relative morphism space
$\pi\colon\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U):=\{(x,g):g\in\operatorname{Hom}(Y_{x},Z),x\in U\}\rightarrow U,\quad(x,g)\mapsto x$
over $U$, of which connected components are complex analytic spaces by Douady’s theorem [2]. Let $\sigma\colon U\rightarrow\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U)$ be the continuous section defined by $\sigma(x):=(x,f_{x})$, and denote by $\mathcal{H}$ the connected component of $\operatorname{Hom}_{U}(\phi^{-1}(U),Z\times U)$ containing $\sigma(U)$.
For each integer $k\geq 1$, we define a holomorphic mapping $\Phi_{k}\colon\mathcal{H}\rightarrow(Z\times\cdots\times Z)\times U$ by
$\Phi_{k}(x,g):=(g(s_{1}(x)),\dots,g(s_{k}(x)),x),$
and define
$A_{k}:=\Phi_{k}^{-1}(\{(f(s_{1}(x)),\dots,f(s_{k}(x)),x):x\in U\}).$
Then $\{A_{k}\}_{k\geq 1}$ is a descending chain of complex analytic subsets in $\mathcal{H}$. By the analytic continuation principle, it holds that
$\bigcap_{k=1}^{\infty}A_{k}=\sigma(U).$
Because the local rings of $\mathcal{H}$ are Noetherian, the descending chain $\{A_{k}\}_{k\geq 1}$ must stabilize in a neighborhood of any point in $\sigma(U)$. Consequently, the restriction $\pi|_{\sigma(U)}$ is a bijective holomorphic mapping onto its image. Since $U$ is smooth, Zariski’s main theorem [4] immediately yields that $\sigma$ is holomorphic. By the universal property of the relative morphism space, the mapping $f|_{\phi^{-1}(U)}$ itself is holomorphic.
Since $f$ is continuous and the singular locus $Y-\phi^{-1}(U)$ is a proper analytic subset of $Y$, Riemann’s extension theorem applied locally in the coordinate charts of $Z$ guarantees that $f$ is holomorphic on the entirety of $Y$. ∎
The application of Lemma 2.1 converts algebraic data, namely the stabilization of descending analytic chains, into global analytic rigidity. We now apply this machinery to elevate continuous homomorphisms between elliptic fibrations and Abelian schemes into holomorphic morphisms. By utilizing the group law of the elliptic fibration, a single torsion-free holomorphic section generates the required sequence of disjoint sections.
@@
Let $\pi\colon E\rightarrow B$ be a smooth elliptic fibration and $\varphi\colon A\rightarrow B$ an Abelian scheme over a complex manifold $B$. Let $f\colon E\rightarrow A$ be a continuous mapping such that $\pi=\varphi\circ f$ and
$f_{b}\colon\pi^{-1}(b)\rightarrow\varphi^{-1}(b),\quad P\mapsto f(P)$
is a homomorphism for every $b\in B$. Assume there exists a holomorphic section $\sigma$ of $\pi$ such that $f\circ\sigma$ is holomorphic and $\sigma(b)$ is torsion-free for every $b\in B$. Then $f$ is holomorphic.
@@
For each integer $k\geq 1$, we define a section $s_{k}\colon B\rightarrow E$ by $s_{k}(b):=k\sigma(b)$, where the multiplication is taken with respect to the group structure of the elliptic fibration $\pi$. Since $\sigma(b)$ is torsion-free for every $b\in B$, the sequence of sections $\{s_{k}\}_{k\geq 1}$ has mutually disjoint images.
Since $f$ is continuous, we have that $f_{b}$ is also continuous for all $b\in B$. But any continuous homomorphism between complex Lie groups is holomorphic, ensuring that $f|_{\pi^{-1}(b)}$ is holomorphic for all $b\in B$.
Because the restriction $f_{b}\colon\pi^{-1}(b)\rightarrow\varphi^{-1}(b)$ is a group homomorphism for every $b\in B$, we obtain
$(f\circ s_{k})(b)=f_{b}(k\sigma(b))=kf_{b}(\sigma(b))=k(f\circ\sigma)(b)$
for all $b\in B$. Therefore we have the identity $f\circ s_{k}=\mu_{k}\circ(f\circ\sigma)$, where $\mu_{k}\colon A\rightarrow A$ denotes the integer multiplication morphism on the smooth Abelian scheme $A$. Since $A$ is Abelian, the mapping $\mu_{k}$ is holomorphic. By the assumption that $f\circ\sigma$ is holomorphic, the composition $f\circ s_{k}$ is holomorphic for every integer $k\geq 1$.
The desired result then follows immediately from Lemma 2.1. ∎
Remark 2.3.
The proof of Corollary 2.2 encapsulates the overarching theme of this section: rigid global structures in complex geometry can be rigorously enforced by combining proper fibrations with the positive constraint of a torsion-free section. The algebraic stabilization within the relative morphism space serves as the essential bridge between continuous topological data and complex analytic rigidity.
3 Rigidity of Mappings from Kobayashi Hyperbolic Manifolds
@@
Step 1: Since $H\in\operatorname{div}(X)$ is ample, the complex manifold $X$ is projective by Kodaira’s embedding theorem [3]. Let
@@
be the punctured Riemann sphere consisting of the regular values of $f$. For any $t\in\Sigma$, the fiber $X^{t}:=f^{-1}(t)$ is a smooth projective subvariety of $X$ of pure dimension n1n-1. Since $\phi^{t}:=\phi|_{X^{t}}$ is holomorphic and $Y$ is projective, Remmert’s proper mapping [4] theorem and Chow’s theorem [3] ensure that the image $Y^{t}:=\phi(X^{t})$ is a projective subvariety of $Y$. We first establish the injectivity of the restriction $\phi^{t}$ for any fixed $t\in\Sigma$. For the sake of contradiction, we now assume that $\phi^{t}$ is not injective.
@@
Suppose now that $\dim(Y^{t})=n-2$. Take a generic point $y\in Y^{t}$. Then the fiber $C$ of $\phi^{t}$ over $y$ is a projective curve. Since $H\in\operatorname{div}(X)$ is a very ample divisor, $H^{t}:=X^{t}\cap H$ is very ample in $X^{t}$, making $H^{t}\cap C$ non-empty and $\phi(H^{t})=Y^{t}$. Because $\phi|_{H}$ is injective by assumption, the mapping degree of $\phi^{t}|_{H^{t}}$ onto its image is exactly 1. By choosing $y$ to be generic, we ensure that the restriction $\phi^{t}|_{H^{t}}$ is unramified at its unique preimage $x\in H^{t}$, rendering its differential an isomorphism at $x$. As a result, the tangent space $T_{x}H^{t}$ intersects trivially with the kernel of the differential of $\phi^{t}$ at $x$, which is precisely $T_{x}C$. This proves that $C$ intersects $H^{t}$ transversally, and hence their intersection product is $H^{t}\cdot C=1$. Because $H$ is a very ample divisor, intersecting a projective curve $C$ at a single point with a multiplicity of exactly 1 forces $C$ to be rational, that is, bi-holomorphic to $\mathbb{P}^{1}$. However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem [1] on hyperbolicity. This is a contradiction. Therefore, we have in fact $\dim(Y^{t})=n-1$, so the regular mapping $\phi^{t}\colon X^{t}\rightarrow Y$ sends its domain of definition onto a projective variety $Y^{t}$ of equal dimension n1n-1.
@@
Because $\phi^{t}_{2}$ is a surjective morphism with connected fibers between projective varieties of the same dimension, its generic fiber is a single point, implying that $\phi^{t}_{2}$ is a birational morphism. Assume to the contrary that $\phi^{t}$ admits a fiber $F$ of positive dimension. Since $\phi^{t}_{1}$ is finite, this fiber $F$ must be contracted to a point entirely by $\phi^{t}_{2}$. Therefore, the exceptional locus of the birational morphism $\phi^{t}_{2}$ is non-empty. Now, recall from birational geometry that the exceptional locus of a proper birational morphism from a smooth projective variety is uniruled, meaning it is covered by rational curves. However, as $X$ is Kobayashi hyperbolic, by Brody’s theorem [1], it contains no rational curves, which leads to a contradiction. This proves that $\phi^{t}$ is indeed a finite morphism.
@@
for all $t\in\mathbb{P}^{1}$. Since $Y$ is a smooth projective variety and $g\colon Y\rightarrow\mathbb{P}^{1}$ is a holomorphic fibration, Bertini’s theorem [6] ensures that the general fiber of $g$ is smooth, and therefore normal. Thus, for a generic $t\in\Sigma$, the finite morphism $\phi^{t}\colon X^{t}\rightarrow Y^{t}$ of degree 1 onto a normal variety is an isomorphism by Zariski’s main theorem [6]. This in particular proves that $\phi$ is injective almost everywhere.
@@
of the branched covering $\Phi|_{Z}$. Since $\Phi|_{Z}$ admits a continuous section over $\Sigma$, it is in fact not branched over $\Sigma$. Moreover, any continuous section of an unbranched analytic covering is necessarily holomorphic, so $\phi^{*}$ is holomorphic on $\Sigma$. Because $\phi$ is holomorphic along fibers and the section $\phi^{*}$ is also holomorphic, Osgood’s lemma [4] on separate holomorphicity guarantees that $\phi$ is holomorphic on $f^{-1}(\Sigma)$. Since $\phi$ is continuous and the singular locus $X-f^{-1}(\Sigma)$ forms a proper algebraic subset, Riemann’s extension theorem guarantees that $\phi$ is holomorphic on the entirety of $X$.
Having established that $\phi\colon X\rightarrow Y$ is a proper holomorphic mapping of degree 1 between smooth projective varieties, it remains to show it is an isomorphism. But, if $\phi$ were not an isomorphism, as it is injective almost everywhere, it would be a birational map contracting an exceptional divisor to a lower-dimensional subvariety. Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$.
Quod erat demonstrandum. ∎
Remark 3.2.
Theorem 3.1 illustrates another facet of mapping regularity in complex geometry. While the results in Section 2 rely on an infinite sequence of disjoint sections to propagate holomorphicity, this theorem demonstrates that a single very ample hypersurface provides sufficient algebraic anchoring to elevate a continuous mapping to a global bi-holomorphic isomorphism. The intrinsic absence of rational curves in the Kobayashi hyperbolic manifold prevents the formation of exceptional loci, ensuring that the induced birational morphism is necessarily an isomorphism.